# Reliable algorithm selection for machine learning-guided design

### v1

@@ Results
While our method’s error rate was not under $\alpha$ due to density ratio estimation error, it was much lower than those of the prediction-only and CalibratedForecasts methods, which had $100\%$ error rates for most values of $\tau$ (Fig. 4a). Furthermore, when our method selected configurations that were unsuccessful, they still produced mean design labels close to $\tau$ (Fig. 4b). Our method assesses configurations more accurately at the cost of selections, $\hat{\Lambda}$, that are higher variance (Fig. 4b), due to the reduced effective sample size of the weighted labeled data.
@@
with $q=0$ achieved low error rates, often even zero, for $\tau<0.37$, but was very conservative: it stopped selecting configurations for greater $\tau$ (Fig. 4a). Using $q\in\{0.5,1\}$ increased selection rates at the cost of much higher error rates, though, similar to our method, the effects of the errors were not severe for $q=0.5$: the mean design labels of unsuccessful configurations were not far below $\tau$ (Fig. 4b).
@@ 6 Discussion
We introduced an algorithm selection method for machine learning-guided design, which selects design algorithm configurations that will be successful with high probability—that is, produce a distribution of design labels that satisfies a user-specified success criterion. It does so by using held-out labeled data to characterize and then undo how prediction error biases the assessment of whether a configuration is successful. Though the present work focuses on success in a single “round” of design, it can also provide a principled decision-making framework for multi-round design efforts in which a top priority is that the designs at each round achieve certain criteria—for example, to justify resources for acquiring their labels.
As with other uncertainty quantification methods that achieve frequentist-style guarantees under covariate shift [47, 15, 30, 25], the method uses the density ratios between the design and labeled distributions. Advances in density ratio estimation techniques will strengthen the method’s performance in settings where these density ratios are not known—in particular, techniques that are optimized specifically for importance-weighted mean estimation. Looking forward, we encourage continued work on how to address predictive uncertainty in machine learning-guided design with respect to how it directly impacts endpoints or decisions of interest, rather than with general-purpose notions of uncertainty [19].
@@ 7 Acknowledgments
Our gratitude goes to Anastasios N. Angelopoulos, Stephen Bates, Richard Bonneau, Andreas Loukas, Ewa Nowara, Stephen Ra, Samuel Stanton, Nataša Tagasovska, and Tijana Zrnic for helpful discussions and feedback on this work.
@@
Appendix A Algorithms and proofs
Algorithm 3 Prediction-powered p-value for testing $H_{\lambda}:\theta_{\lambda}<\tau$ (finite sample-valid)
Inputs: Predictions for designs, $\{\hat{y}^{\prime}_{i}\}_{i=1}^{N}$; labeled data and their predictions, $\{(x_{j},y_{j},\hat{y}_{j})\}_{j=1}^{n}$; fine grid spacing, $\delta>0$.
@@ Appendix A Algorithms and proofs
1: $w_{j}\leftarrow\textsc{DensityRatio}(x_{j}),\;j=1,\ldots,n$
@@
3:     $L_{\alpha}\leftarrow\textsc{PPMeanLB}(\alpha,\;\{\hat{y}^{\prime}_{i}\}_{i=1}^{N},\;\{(w_{j},y_{j},\hat{y}_{j})\}_{j=1}^{n})$
@@
Inputs: Significance level, $\alpha\in[0,1]$; predictions for designs, $\{\hat{y}^{\prime}_{i}\}_{i=1}^{N}$; labeled data and their density ratios and predictions, $\{(w_{i},y_{i},\hat{y}_{i})\}_{i=1}^{n}$; range, $[L,U]$, of $g(Y)$; bound, $B$, on the density ratios.
@@
1: $\hat{\mu}_{\text{lower}}\leftarrow\textsc{MeanLB}(0.1\cdot\alpha,\;\{g(\hat{y}^{\prime}_{i})\}_{i=1}^{N},\;[L,U])$
2: $\Delta_{\text{lower}}\leftarrow\textsc{MeanLB}(0.9\cdot\alpha,\;\{w_{j}\cdot(g(\hat{y}_{j})-g(y_{j}))\}_{j=1}^{n},\;[B(L-U),B(U-L)])$
3: $L\leftarrow\hat{\mu}_{\text{lower}}+\Delta_{\text{lower}}$
Algorithm 5 MeanLB: Confidence lower bound on a mean (finite sample-valid; Waudby-Smith and Ramdas [50])
@@
4: $M_{0}^{-}(m)\leftarrow 1,m\in\mathcal{A}$
5:for $t=1,\ldots,n$ do
6:      $\hat{\mu}_{t}\leftarrow\frac{0.5+\sum_{j=1}^{t}z_{j}}{t+1},\hat{\sigma}^{2}_{t}\leftarrow\frac{0.25+\sum_{j=1}^{t}(z_{j}-\hat{\mu}_{t})^{2}}{t+1},\lambda_{t}\leftarrow\sqrt{\frac{2\log(2/\alpha)}{n\hat{\sigma}^{2}_{t-1}}}$
7:     for $m\in\mathcal{A}$ do
8:          $M_{t}^{+}(m)\leftarrow\left(1+\min\{\lambda_{t},\frac{0.5}{m}\}(z_{t}-m)\right)M_{t-1}^{+}(m)$
9:          $M_{t}^{-}(m)\leftarrow\left(1-\min\{\lambda_{t},\frac{0.5}{1-m}\}(z_{t}-m)\right)M_{t-1}^{-}(m)$
10:          $M_{t}(m)=0.5\cdot\max\{M_{t}^{+}(m),M_{t}^{-}(m)\}$
11:         if $M_{t}(m)\geq 1/\alpha$ then
12:               $\mathcal{A}\leftarrow\mathcal{A}\setminus\{m\}$
13:         end if
14:     end for
15:end for
16: $B\leftarrow\min\mathcal{A}\cdot(U-L)+L$
Theorem 2.
For any error rate, $\alpha\in[0,1]$, function $g:\mathcal{Y}\rightarrow\mathbb{R}$, and threshold $\tau\in\mathbb{R}$, Algorithm 1 using Algorithm 2 as the p-value subroutine returns a subset of configurations from the menu, $\hat{\Lambda}\subseteq\Lambda$, that satisfies
@@
where $\theta_{\lambda}\equiv\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$, $n$ and $N$ are the amounts of labeled data and designs sampled per configuration, respectively, $\frac{n}{N}\rightarrow r$ for some $r\in(0,1)$, and the probability is over random draws of the labeled data and the designs for all configurations on the menu.
The proofs of Theorems 1 and 2 rely on establishing the validity of the p-values computed by Algorithms 3 and 2, respectively. Both algorithms use the framework of prediction-powered inference [2], which estimates a population-level estimand of interest—here, $\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$—by combining predictions for $Y$ with an estimand-specific characterization of prediction error called the rectifier. For our estimand, these two components emerge from the following simple decomposition:
@@
where the first summand is the mean prediction for designs, and the second is the mean prediction bias, which serves as the rectifier. More generally, prediction-powered inference encompasses any estimand that can be expressed as the minimizer of an expected convex loss (see Appendix B.1), in which case the rectifier can be derived from a similar decomposition of that loss’s gradient.
To estimate the two terms in Eq. 4, we can use the predictions for $N$ designs to estimate the first, and the $n$ labeled instances to estimate the second, the rectifier. In particular, since the rectifier is an expectation over designs and their labels, we leverage the covariate shift relationship between the distributions of the designs and the labeled data to rewrite it as
@@
This last expression can be estimated using the labeled data, where each instance is weighted by the density ratio between the design and label distributions. Adding together the estimates of the two terms in Eq. 4 yields a prediction-powered estimate of $\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$, and a prediction-powered p-value can be computed based on this estimate. Specifically, Algorithm 2 follows standard protocol for constructing asymptotically valid p-values, by first deriving the asymptotic null distribution of this prediction-powered estimate, and then evaluating its survival function to produce a p-value. Algorithm 3 relies on inverting finite sample-valid, prediction-powered confidence lower bounds on $\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$ to obtain p-values.
We now briefly describe why the prediction-powered estimate of $\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$ is statistically more powerful—and yields correspondingly more powerful p-values—than ignoring the $N$ predictions and using only the $n$ weighted labeled instances to estimate it, rewritten as
@@
$=\mathbb{E}_{X,Y\sim P_{\text{lab}}\cdot P_{Y\mid X}}\left[\frac{p_{X;\lambda}(X)}{p_{\text{lab}}(X)}g(Y)\right].$ (5)
Critically, we assume that $N\gg n$, as it is typically far cheaper to computationally generate designs than it is to acquire labeled data. For $N\gg n$, the variance of the estimate of the first term in Eq. 4 will be negligible compared to that of the rectifier; indeed, we can make it arbitrarily small by computationally generating more designs from a given configuration. The variance of the prediction-powered estimate is therefore dominated by the variance of the rectifier estimate using the $n$ labeled instances. Furthermore, the rectifier will be small so long as the predictions carry some information about the labels, so the variance of its estimate will likely be well below that of an estimate of Eq. A using those same $n$ labeled instances.
A.1 Proof of Theorem 2 (asymptotically valid design algorithm selection)
@@
We first derive the asymptotic distribution of $\hat{\theta}=\hat{\mu}_{\hat{y}^{\prime}}+\hat{\Delta}$, where $\hat{\mu}_{\hat{y}^{\prime}}=(1/N)\sum_{i=1}^{N}g(\hat{y}^{\prime}_{i})$ is the average prediction for $N$ designs, and $\hat{\Delta}=(1/n)\sum_{j=1}^{n}w_{j}(g(y_{j})-g(\hat{y}_{j}))$ where $w_{j}=p_{X;\lambda}(x_{j})/p_{\text{lab}}(x_{j})$ is the importance-weighted average bias of predictions for the $n$ labeled instances. The central limit theorem implies that
@@
where $\sigma^{2}_{\hat{y}^{\prime}}=\text{Var}_{X\sim P_{X;\lambda}}[g(f(X))]$ and $\sigma^{2}_{y^{\prime}-\hat{y}^{\prime}}=\text{Var}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[g(Y)-g(f(X))]$. Since $\hat{\mu}_{\hat{y}^{\prime}}$ and $\hat{\Delta}$ are independent, we can apply the continuous mapping theorem to their sum to get that
$\sqrt{N}(\hat{\mu}_{\hat{y}^{\prime}}+\hat{\Delta}-\mathbb{E}[\hat{\mu}_{\hat{y}^{\prime}}+\hat{\Delta}])$ $=\sqrt{N}(\hat{\mu}_{\hat{y}^{\prime}}-\mathbb{E}[\hat{\mu}_{\hat{y}^{\prime}}])+\sqrt{\frac{N}{n}}\sqrt{n}(\hat{\Delta}-\mathbb{E}[\hat{\Delta}])$
@@
where recall $n/N\rightarrow r$. Since
$\mathbb{E}[\hat{\mu}_{\hat{y}^{\prime}}+\hat{\Delta}]=\mathbb{E}_{X\sim P_{X;\lambda}}[g(f(X))]+\mathbb{E}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[g(Y)-g(f(X))]=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]:=\theta_{\lambda},$
@@
where $\hat{\sigma}^{2}$ is any consistent estimate of $\sigma^{2}_{\hat{y}^{\prime}}+\frac{1}{r}\sigma^{2}_{\hat{y}^{\prime}-y^{\prime}}$, satisfies Eq. 6. We use the estimate $\hat{\sigma}^{2}=\hat{\sigma}^{2}_{\hat{y}^{\prime}}+\frac{N}{n}\hat{\sigma}^{2}_{\hat{y}-y}$ where $\hat{\sigma}^{2}_{\hat{y}^{\prime}}=\frac{1}{N}\sum_{i=1}^{N}(g(\hat{y}^{\prime}_{i})-\hat{\mu}_{\hat{y}^{\prime}})^{2}$ and $\hat{\sigma}^{2}_{\hat{y}-y}=\frac{1}{n}\sum_{j=1}^{n}(w_{j}[g(y_{j})-g(\hat{y}_{j})]-\hat{\Delta})^{2}$, which is consistent since $\hat{\sigma}^{2}_{\hat{y}^{\prime}}$ and $\hat{\sigma}^{2}_{\hat{y}-y}$ are consistent estimates of $\sigma^{2}_{\hat{y}^{\prime}}$ and $\sigma^{2}_{\hat{y}^{\prime}-y^{\prime}}$, respectively.
@@
$\limsup_{n,N\rightarrow\infty}\text{FWER}:=\limsup_{n,N\rightarrow\infty}\mathbb{P}\left(\bigcup_{\lambda\in\Lambda:\;\theta_{\lambda}<\tau}\lambda\in\hat{\Lambda}\right)$ $\leq\sum_{\lambda\in\Lambda:\;\theta_{\lambda}<\tau}\limsup_{n,N\rightarrow\infty}\mathbb{P}\left(\lambda\in\hat{\Lambda}\right)$ (7)
$=\sum_{\lambda\in\Lambda:\;\theta_{\lambda}<\tau}\limsup_{n,N\rightarrow\infty}\mathbb{P}\left(p_{\lambda}\leq\frac{\alpha}{|\Lambda|}\right)$ (8)
|$\leq\left|\{\lambda\in\Lambda:\;\theta_{\lambda}<\tau\}\right|\cdot\frac{\alpha}{|\Lambda|}$ (9)
$\leq|\Lambda|\cdot\frac{\alpha}{|\Lambda|}=\alpha,$ (10)
@@
A.2 Proof of Theorem 1 (finite sample-valid design algorithm selection)
@@
We first show that the p-value computed by Algorithm 3, $p_{\lambda}$, has finite-sample validity, meaning that under the null hypothesis, $H_{\lambda}:\theta_{\lambda}<\tau$, we have $\mathbb{P}(p_{\lambda}\leq u)\leq u,\forall u\in[0,1]$. First, the confidence lower bound, $L_{\alpha}$, computed by PPMeanLB (Alg. 4) is valid: it satisfies $\mathbb{P}(\theta_{\lambda}\geq L)\geq 1-\alpha$. This follows from the fact that MeanLB produces valid confidence lower bounds, $\hat{\mu}_{\text{lower}}$ and $\Delta_{\text{lower}}$, for $\mathbb{E}_{X\sim P_{X;\lambda}}[g(f(X))]$ and $\mathbb{E}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[g(Y)-g(f(X))]$, respectively (Theorem 3 from Waudby-Smith and Ramdas [50]). Adding together these bounds, following the decomposition in Eq. 4, therefore yields a valid confidence lower bound, $L_{\alpha}$, for $\theta_{\lambda}$. Algorithm 3 then constructs a p-value by inverting $L_{\alpha}$:
$p_{\lambda}=\inf\{\alpha\in[0,1]:L_{\alpha}>\tau\}.$ (11)
@@
$\mathbb{P}(p_{\lambda}\leq u)\leq\mathbb{P}(\theta_{\lambda}<L_{u})=1-\mathbb{P}(\theta_{\lambda}\geq L_{u})=1-(1-u)=u,$
where the first inequality follows from the definition of $p_{\lambda}$ in Eq. 11 and the fact that $\theta_{\lambda}<\tau$, and the second equality comes from the validity of $L_{u}$.

### v2

@@ Results
The prediction-only and CalibratedForecasts methods had $100\%$ error rates for much of the range of $\tau$ considered. Our method had much lower error rates, though greater than $\alpha$ for $\tau>0.32$ due to density ratio estimation error (Fig. 4a, top). When our method selected configurations that were unsuccessful, their mean design labels were still close to $\tau$ (Fig. 4b; the shaded green region does not extend far below the dashed diagonal line). Our method assesses configurations more accurately at the cost of selected subsets, $\hat{\Lambda}$, that are higher variance (Fig. 4b; the shaded green region is wider than other shaded regions), due to the reduced effective sample size of the weighted labeled data.
@@
with $q=0$ had low error rates, even zero, for $\tau<0.37$, but was the most conservative method: it stopped selecting any configurations for greater $\tau$ (Fig. 4a, bottom). Using $q\in\{0.5,1\}$ maintained $100\%$ selection rates for broader ranges of $\tau$ at the cost of much higher error rates. However, similar to our method, the errors’ consequences were not severe, especially for $q=0.5$: the mean design labels of unsuccessful configurations were not far below $\tau$ (Fig. 4b).
Overall, compared to any alternative method except GMMForecasts with $q=0$, our method had lower error rates over the range of $\tau$ for which the alternative had non-zero selection rates (Fig. 4a). Our method also maintained non-zero selection rates for a broader range of $\tau$ than GMMForecasts with any $q$. These results—which hold even with estimated density ratios, and holding out labeled data—illustrate the benefits of quantifying the consequences of prediction error on downstream quantities or decisions of interest, over focusing on the uncertainties of individual predictions.
@@ 6 Discussion
We introduced an algorithm selection method for machine learning-guided design, which selects design algorithm configurations that will be successful with high probability—that is, produce a distribution of design labels that satisfies a user-specified, population-level success criterion. It does so by using held-out labeled data to characterize and then undo how prediction error biases the assessment of whether a configuration is successful, leveraging techniques from prediction-powered inference (Angelopoulos et al.,, 2023). Though the present work focuses on success in a single “round” of design, it could also assist multi-round efforts in which a top priority is that designs at each round achieve certain criteria—for example, to justify resources for acquiring their labels.
As with other uncertainty quantification methods that achieve frequentist-style guarantees under covariate shift (Tibshirani et al.,, 2019; Fannjiang et al.,, 2022; Prinster et al.,, 2023; Jin and Candès,, 2023), the method uses the density ratios between the design and labeled distributions. Advances in density ratio estimation techniques may strengthen the method’s performance in settings where these density ratios are not known—in particular, techniques that are tailored for importance-weighted mean estimation. Another promising direction is the incorporation of multiple testing procedures that respect structure among configurations on the menu, which may enable less conservative selection than the Bonferroni correction. Looking forward, we encourage continued work on how to address predictive uncertainty in machine learning-guided design with respect to how it directly impacts endpoints or decisions of interest, rather than with general-purpose notions of uncertainty (Greenman et al.,, 2025).
@@ 7 Acknowledgments
Our gratitude goes to Anastasios N. Angelopoulos, Stephen Bates, Richard Bonneau, Kyunghyun Cho, Andreas Loukas, Ewa Nowara, Stephen Ra, Samuel Stanton, Nataša Tagasovska, and Tijana Zrnic for helpful discussions and feedback on this work.
@@
Appendix A Additional Algorithms and Proofs
Algorithm 3 Prediction-powered p-value for testing $H_{\lambda}:\theta_{\lambda}:=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]<\tau$ (finite sample-valid)
Inputs: Predictions for designs, $\{\hat{y}^{\lambda}_{i}\}_{i=1}^{N}$; labeled data and their predictions, $\{(x_{j},y_{j},\hat{y}_{j})\}_{j=1}^{n}$; small grid spacing, $\delta>0$.
@@ Appendix A Additional Algorithms and Proofs
1: $w_{j}\leftarrow\textsc{DensityRatio}(x_{j}),\;j=1,\ldots,n$triangleright $\text{{DensityRatio}}(\cdot):=p_{X;\lambda}(\cdot)/p_{\text{lab}}(\cdot)$
@@
3:     $L_{\alpha}\leftarrow\textsc{PPMeanLB}(\alpha,\;\{\hat{y}^{\lambda}_{i}\}_{i=1}^{N},\;\{(y_{j},\hat{y}_{j},w_{j})\}_{j=1}^{n})$
@@
Inputs: Significance level, $\alpha\in[0,1]$; predictions for designs, $\{\hat{y}^{\lambda}_{i}\}_{i=1}^{N}$; labels, predictions, and density ratios for labeled data, $\{(y_{i},\hat{y}_{i},w_{i})\}_{i=1}^{n}$; range, $[L,U]$, of $g(Y)$; bound, $D$, on the density ratios.
@@
1: $\hat{\mu}_{\text{lower}}\leftarrow\textsc{MeanLB}(0.1\cdot\alpha,\;\{g(\hat{y}^{\lambda}_{i})\}_{i=1}^{N},\;[L,U])$
2: $\hat{\Delta}_{\text{lower}}\leftarrow\textsc{MeanLB}(0.9\cdot\alpha,\;\{w_{j}\cdot(g(y_{j})-g(\hat{y}_{j}))\}_{j=1}^{n},\;[D(L-U),D(U-L)])$
3: $L\leftarrow\hat{\mu}_{\text{lower}}+\hat{\Delta}_{\text{lower}}$
Algorithm 5 MeanLB: Confidence lower bound on a mean (finite sample-valid; due to Waudby-Smith & Ramdas (2023) (Waudby-Smith and Ramdas,, 2023))
@@
4:for $t=1,\ldots,n$ do
5:      $\hat{\mu}_{t}\leftarrow\frac{0.5+\sum_{i=1}^{t}z_{i}}{t+1},\hat{\sigma}^{2}_{t}\leftarrow\frac{0.25+\sum_{i=1}^{t}(z_{i}-\hat{\mu}_{t})^{2}}{t+1},\lambda_{t}\leftarrow\sqrt{\frac{2\log(2/\alpha)}{n\hat{\sigma}^{2}_{t-1}}}$
6:     for $m\in\mathcal{A}$ do
7:          $M_{t}^{+}(m)\leftarrow\left(1+\min\{\lambda_{t},\frac{0.5}{m}\}(z_{t}-m)\right)M_{t-1}^{+}(m)$
8:         if $M_{t}^{+}(m)\geq 1/\alpha$ then
9:               $\mathcal{A}\leftarrow\mathcal{A}\setminus\{m\}$
10:         end if
11:     end for
12:end for
13: $B\leftarrow\min\mathcal{A}\cdot(U-L)+L$
Theorem A.1.
For any error rate $\alpha\in[0,1]$, function $g:\mathcal{Y}\rightarrow\mathbb{R}$, and threshold $\tau\in\mathbb{R}$, Algorithm 1 using Algorithm 2 as the p-value subroutine returns a subset of configurations from the menu, $\hat{\Lambda}\subseteq\Lambda$, that satisfies
@@
where $\theta_{\lambda}:=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$, $n$ and $N$ are the amounts of labeled data and designs from each configuration, respectively, $\frac{n}{N}\rightarrow r$ for some $r\in(0,1)$, and the probability is over random draws of labeled data and designs from all configurations on the menu.
A.1 Proofs of Theorems 3.1 and A.1
The proofs of Theorems 3.1 and A.1 rely on establishing the validity of the p-values computed by Algorithms 3 and 2, respectively. Both algorithms use the framework of prediction-powered inference (Angelopoulos et al.,, 2023), which constructs confidence intervals and p-values for an estimand of interest—here, the population-level metric, $\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$—by combining predictions for $Y$ with an estimand-specific characterization of prediction error called the rectifier. For our estimand, these two components emerge from the following simple decomposition of the population-level metric:
@@
where $f$ is any predictive model. The first summand is the mean prediction for designs, and the second is the mean prediction bias, which serves as the rectifier. More generally, prediction-powered inference encompasses any estimand that can be expressed as the minimizer of an expected convex loss (see Appendix B.1), in which case the rectifier can be derived from a similar decomposition of the expected loss gradient.
Let $\{x^{\lambda}_{i}\}_{i=1}^{N}$ denote $N$ designs from configuration $\lambda$, and let $\{(x_{j},y_{j})\}_{j=1}^{n}$ denote the $n$ labeled instances. We can use the predictions for the designs to estimate the first summand in Eq. 4 as
$\hat{\mu}:=\frac{1}{N}\sum_{i=1}^{N}g(f(x^{\lambda}_{i})),$ (5)
and use the $n$ labeled instances to estimate the second, the rectifier. Specifically, we leverage the covariate shift relationship between the design and labeled data to rewrite the rectifier as
@@
This final expression can be estimated using the labeled data, where each instance is weighted by the density ratio between the design and label distributions:
$\hat{\Delta}:=\frac{1}{n}\sum_{j=1}^{n}\frac{p_{X;\lambda}(x_{j})}{p_{\text{lab}}(x_{j})}\left(g(y_{j})-g(f(x_{j})\right).$ (6)
Adding together the two estimates in Eqs. 5 and 6 yields a prediction-powered estimate of the population-level metric,
$\hat{\theta}:=\hat{\mu}+\hat{\Delta},$ (7)
from which a prediction-powered p-value can be computed. Specifically, Algorithm 2 follows standard protocol for constructing asymptotically valid p-values, by first deriving the asymptotic null distribution of the prediction-powered estimate, and then evaluating its survival function to produce a p-value. Alternatively, Algorithm 3 inverts finite sample-valid, prediction-powered confidence lower bounds on the population-level metric to obtain finite sample-valid p-values.
We now briefly describe why the prediction-powered estimate of the population-level metric, Eq. 7, generally has lower variance—and yields correspondingly more powerful p-values—than ignoring the $N$ predictions and using only the $n$ weighted labeled instances to estimate the population-level metric. Specifically, the covariate shift relationship between the design and labeled data allows us to rewrite the population-level metric as
@@
$=\mathbb{E}_{X,Y\sim P_{\text{lab}}\cdot P_{Y\mid X}}\left[\frac{p_{X;\lambda}(X)}{p_{\text{lab}}(X)}g(Y)\right],$ (8)
and correspondingly construct a “labeled-only” estimate,
$\frac{1}{n}\sum_{j=1}^{n}\frac{p_{X;\lambda}(x_{j})}{p_{\text{lab}}(x_{j})}g(y_{j}).$ (9)
We now compare the asymptotic variances of the prediction-powered estimate, Eq. 7, and the labeled-only estimate, Eq. 9. Critically, we assume that $N\gg n$, as it is typically far cheaper to generate designs in silico than it is to acquire labeled data. The central limit theorem gives the asymptotic variance of the prediction-powered estimate as
$\frac{1}{N}\text{Var}(g(f(X^{\lambda})))+\frac{1}{n}\text{Var}\left(\frac{p_{X;\lambda}(X)}{p_{\text{lab}}(X)}\left(g(Y)-g(f(X))\right)\right).$
For $N\gg n$, this expression is dominated by the second term, the variance of the weighted prediction error scaled by $1/n$ (indeed, we can make the first term arbitrarily small by computationally generating more designs). The asymptotic variance of the labeled-only estimate is the variance of the weighted labels scaled by $1/n$,
$\frac{1}{n}\text{Var}\left(\frac{p_{X;\lambda}(X)}{p_{\text{lab}}(X)}g(Y)\right).$
So long as the variance of the weighted prediction error is smaller than that of the weighted labels—that is, the predictions explain some of the variance of labels—the asymptotic variance of the prediction-powered estimate will be smaller than that of the labeled-only estimate. That is, the prediction-powered estimate uses information from the predictions to increase its effective sample size compared to the labeled-only estimate.
Proof of Theorem A.1 (asymptotically valid design algorithm selection)
@@
We first derive the asymptotic distribution of $\hat{\theta}:=\hat{\mu}+\hat{\Delta}$, where $\hat{\mu}:=(1/N)\sum_{i=1}^{N}g(\hat{y}^{\lambda}_{i})$ is the average prediction for $N$ designs, and $\hat{\Delta}:=(1/n)\sum_{j=1}^{n}w_{j}(g(y_{j})-g(\hat{y}_{j}))$ where $w_{j}=p_{X;\lambda}(x_{j})/p_{\text{lab}}(x_{j})$ is the importance-weighted average prediction bias for the $n$ labeled instances. The central limit theorem implies that
@@
where $\sigma^{2}_{\text{pred}}:=\text{Var}_{X\sim P_{X;\lambda}}[g(f(X))]$ and $\sigma^{2}_{\text{err}}:=\text{Var}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[g(Y)-g(f(X))]$. Applying the continuous mapping theorem to the sum of $\hat{\mu}$ and $\hat{\Delta}$ yields that
$\sqrt{N}(\hat{\mu}+\hat{\Delta}-\mathbb{E}[\hat{\mu}+\hat{\Delta}])$ $=\sqrt{N}(\hat{\mu}-\mathbb{E}[\hat{\mu}])+\sqrt{\frac{N}{n}}\sqrt{n}(\hat{\Delta}-\mathbb{E}[\hat{\Delta}])$
@@
where recall that $n/N\rightarrow r$. Since
$\mathbb{E}[\hat{\mu}+\hat{\Delta}]=\mathbb{E}_{X\sim P_{X;\lambda}}[g(f(X))]+\mathbb{E}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[g(Y)-g(f(X))]=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]:=\theta_{\lambda},$
@@
where $\hat{\sigma}^{2}$ is any consistent estimate of $\sigma^{2}_{\text{pred}}+\frac{1}{r}\sigma^{2}_{\text{err}}$, satisfies Eq. 10. We use the estimate $\hat{\sigma}^{2}=\hat{\sigma}^{2}_{\text{pred}}+\frac{N}{n}\hat{\sigma}^{2}_{\text{err}}$ where $\hat{\sigma}^{2}_{\text{pred}}:=\frac{1}{N}\sum_{i=1}^{N}(g(\hat{y}^{\lambda}_{i})-\hat{\mu})^{2}$ and $\hat{\sigma}^{2}_{\text{err}}:=\frac{1}{n}\sum_{j=1}^{n}(w_{j}[g(y_{j})-g(\hat{y}_{j})]-\hat{\Delta})^{2}$, which is consistent as $\hat{\sigma}^{2}_{\text{pred}}$ and $\hat{\sigma}^{2}_{\text{err}}$ are consistent estimates of $\sigma^{2}_{\text{pred}}$ and $\sigma^{2}_{\text{err}}$, respectively.
@@
$\limsup_{n,N\rightarrow\infty}\text{FWER}:=\limsup_{n,N\rightarrow\infty}\mathbb{P}\left(\bigcup_{\lambda\in\Lambda:\;\theta_{\lambda}<\tau}\lambda\in\hat{\Lambda}\right)$ $\leq\sum_{\lambda\in\Lambda:\;\theta_{\lambda}<\tau}\limsup_{n,N\rightarrow\infty}\mathbb{P}\left(\lambda\in\hat{\Lambda}\right)$
$=\sum_{\lambda\in\Lambda:\;\theta_{\lambda}<\tau}\limsup_{n,N\rightarrow\infty}\mathbb{P}\left(p_{\lambda}\leq\frac{\alpha}{|\Lambda|}\right)$
$\leq|\Lambda|\cdot\frac{\alpha}{|\Lambda|}=\alpha,$
@@
Proof of Theorem 3.1 (finite sample-valid design algorithm selection)
@@
We first show that the p-value computed by Algorithm 3, $p_{\lambda}$, has finite-sample validity, meaning that under the null hypothesis, $H_{\lambda}:\theta_{\lambda}<\tau$, we have $\mathbb{P}(p_{\lambda}\leq u)\leq u,\forall u\in[0,1]$. First, the confidence lower bound, $L_{\alpha}$, computed by PPMeanLB (Alg. 4) is valid: it satisfies $\mathbb{P}(\theta_{\lambda}\geq L_{\alpha})\geq 1-\alpha$. This follows from the fact that MeanLB produces valid confidence lower bounds, $\hat{\mu}_{\text{lower}}$ and $\hat{\Delta}_{\text{lower}}$, for $\mathbb{E}_{X\sim P_{X;\lambda}}[g(f(X))]$ and $\mathbb{E}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[g(Y)-g(f(X))]$, respectively (Theorem 3 from Waudby-Smith and Ramdas, (2023)). Adding together these bounds therefore yields a valid confidence lower bound, $L_{\alpha}$, for $\theta_{\lambda}$. Algorithm 3 then constructs a p-value by inverting $L_{\alpha}$:
$p_{\lambda}=\inf\{\alpha\in[0,1]:L_{\alpha}\geq\tau\}.$ (11)
@@
$\mathbb{P}(p_{\lambda}\leq u)\leq\mathbb{P}(\theta_{\lambda}<L_{u})=1-\mathbb{P}(\theta_{\lambda}\geq L_{u})\leq 1-(1-u)=u,$
where the first inequality follows from the definition of $p_{\lambda}$ in Eq. 11 and the fact that $\theta_{\lambda}<\tau$, and the second inequality comes from the validity of $L_{u}$.